OpenStudy (abhisar):

The electric field, at a distance of 20 cm from the center of a dielectric sphere of radius 10 cm is 100 N/C. Then Electric field intensity at 3 cm distance from the center of sphere is ?

OpenStudy (abhisar):

@vishweshshrimali5 @Vincent-Lyon.Fr

OpenStudy (abhisar):

@Abmon98

OpenStudy (vishweshshrimali5):

Okay lets see how much I remember ;)

OpenStudy (abhisar):

;)

OpenStudy (vishweshshrimali5):

You would be knowing the formula for electric field due to a charged sphere ?

OpenStudy (abhisar):

yes

OpenStudy (abhisar):

I am using gauss law

OpenStudy (vishweshshrimali5):

Great. You can use Gauss' law to find out the formula for electric field

OpenStudy (vishweshshrimali5):

\[\large{E = k\cfrac{Q}{r^2}}\]

OpenStudy (vishweshshrimali5):

I am not using the vectorial form

OpenStudy (abhisar):

that's ok

OpenStudy (abmon98):

i am sorry i have no idea @Abhisar

OpenStudy (abhisar):

It's ok @Abmon98 ! Thanx for responding :)

OpenStudy (vishweshshrimali5):

I used that formula because distance of separation > radius of sphere

OpenStudy (vishweshshrimali5):

Here: r = 20 cm = 0.2 m E = 100 N/C Using this we can find out the value k

OpenStudy (vishweshshrimali5):

* the value of kQ

OpenStudy (abhisar):

I got it :D

OpenStudy (vishweshshrimali5):

Now for the next part since distance of separation < radius of sphere Thus the formula would be: \[\large{E = \cfrac{kQr}{R^3}}\]

OpenStudy (abhisar):

what is the answer you getting btw ?

OpenStudy (vishweshshrimali5):

Will have to calculate it ... Just a minute :)

OpenStudy (abhisar):

yeah sure !

OpenStudy (vishweshshrimali5):

120 N/C ?

OpenStudy (abhisar):

lol..i am getting 1200 N/c....must be a mistake in conversion

OpenStudy (abhisar):

yes it's 120 N/C .. ..... Thanx @vishweshshrimali5 !

OpenStudy (vishweshshrimali5):

yw :)

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (vishweshshrimali5):

Great !